The Laiterie de Coaticook in the Eastern Townships of Quebec produces several types of cheddar cheese. It markets this cheese in four varieties: aged 2 months, 9 months, 15 months, and 2 years. At the producer's store, 2 kg of each variety sells for the following prices: $ 6.00, $ 9.50, $ 11.00, and $ 12.50, respectively. Consider the cheese maker's decision whether to continue to age a particular 2-pound block of cheese. At 2 months, he can either sell the cheese immediately or let it age further. If he sells it now, he will receive $ 6.00 immediately. If he ages the cheese, he must give up the $ 6.00 today to receive a higher amount in the future. What is the IRR (expressed in percent per month) of the investment of giving up $ 60.00 today by choosing to store 20 kg of cheese that is currently 2 months old and instead selling 10 kg of this cheese when it has aged 9 months, 6 kg when it has aged 15 months, and the remaining 4 kg when it has aged 2 years?

The IRR is __ % per month

please solve this question using a financial calculator, not excel :)
Answer #1

Given we are at time T0= 2 month

CF0 = $(60) (initial Cash inflow)

CF1 = $0 (for no selling for time T1 to T6)

F01= 6 (as the first 10 kg of cheese was sold after 9 months ie at T7 = 9 months)

CF2 = 10*9.5 = $95 ( 10 kg multiply by the per kg price i.e $9.5)

F02 =1

CF3 = $0

F03 = 5 for months (T8 to T14)

CF4 = 6*11 = $66

F04 =1

CF5 = $0

F05 = 4 for T16 to T19

CF6 = 4*12.5 = $50

now press CPT than IRR

IRR = 12.88%